### v1

# Quantum droplets in rapidly rotating two-dimensional Bose-Einstein condensates

Abstract
Recent experiments demonstrate that rapidly rotating Bose-Einstein condensates (BECs) near the lowest Landau level can self-organize into interaction-driven persistent droplet arrays. Inspired by this discovery, we investigate the formation and dynamics of single droplet and droplet arrays in rapidly rotating BECs. Guided by a rigorous theorem on localized many-body states for 2D interacting systems in a magnetic field, we construct single droplet and droplet arrays states which are shown to be stationary solutions to the Gross-Pitaevskii equation in the rotating frame. The single droplet is shown to be dynamically stable, which underpins its role as the basic unit in a droplet array. The stability of the droplet arrays is demonstrated by their dynamic formation from a phase engineered initial condensate. Our study sheds light onto the nature of the droplet state in a rapidly rotating BEC and offers a new approach for generating and manipulating quantum droplet arrays through designing the initial condensate phase.
Introduction.—Rotating Bose-Einstein condensates (BECs) have been subjects of intense interest even since the first realizations of the atomic BEC [1, 2]. Earlier focus has been on the creation of quantized vortices in these systems since they are one of the most definitive signatures of superfludity [3, 4, 5, 6, 7]. As experimental control of these systems progresses, they have emerged as a crucial platform for studying topological quantum phases and quantum fluid dynamics[8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. In particular, under the condition of rapid rotation, the atomic gases in the rotating frame can simulate those in a strong magnetic field, thus allowing for the study of quantum Hall physics in bosonic systems [19]. Indeed, various theories have predicted that as the filling fraction decreases the rotating BEC may transition from the vortex lattice phase in the mean-field quantum Hall regime to various strongly correlated phases in the fractional quantum Hall regime [20, 21, 22, 23, 24, 25].
A critical component of reaching the quantum Hall regimes experimentally is the ability to prepare the atoms in the lowest Landau level (LLL), where the kinetic energy is quenched and the interactions dominate. Although signatures of LLL physics have been previously observed by spinning the BEC up close to the centrifugal limit [26, 27], recent experiments have developed a powerful geometric squeezing method that can prepare the rotating BEC in a single Laudau gauge wave function in the LLL [28, 29, 30]. Moreover, these later experiments were able to achieve filling fractions that are an order of magnitude lower compared to those in earlier ones, placing the system firmly in the mean-field quantum Hall regime. Interestingly, this Laudau gauge BEC does not evolve into a vortex lattice but instead spontaneously form 1D persistent droplet arrays[29], characterized by density modulations at the magnetic length scale and vortex streets between the droplets. Although this crystallization instability has been explained by the condensation of magneto-rotons, the underlying mechanism for the droplet formation and its inherent stability are still open questions.
In this work, we develop a theoretical framework to understand the nature of droplet states in rapidly rotating BECs, by combining rigorous analytic insights with extensive numerical simulations of the Gross-Pitaevskii equation (GPE). We first prove a theorem for 2D interacting systems in a uniform magnetic field, showing that they host many-body eigenstates that can localize around any position in the 2D plane. In the context of rapidly rotating BEC, we use this theorem to construct droplet states, i.e., self-bound states that are dynamically stable [31, 32, 33]. The linear stability of the droplet is explicitly demonstrated by solving the Bogoliubov-de Gennes equations in the rotating frame. Furthermore, we devise a method to construct stationary droplet arrays of any configuration using the single droplet state as the unit. To show that such arrays are stable structures, we adopt a phase-engineering approach and achieve controlled dynamic formation of 2D droplet lattice and circular droplet array. Finally, we show that such a approach can even create droplet arrays in which the droplets perform cyclotron motion. Our work places the concept of quantum droplet state in a rotating BEC on a firm theoretical ground and establishes a novel approach for generating and manipulating droplet arrays, which may provide a pathway to exploring correlated quantum dynamics in these systems.
Localized eigenstates theorem.— We begin by proving a theorem for 2D interacting systems in a magnetic field which will form the basis of our discussion on quantum droplet states in such systems. We show that in these systems i) there exists a class of localized many-body eigenstates that are related to each other by the magnetic translation of the center of mass degree of freedom and ii) these localized states can be set in motion as a whole and form a class of dynamical many-body states for which the probability density performs cyclotron motion without change. The Hamiltonian of a 2D interacting system in a magnetic field reads
$\hat{H}=\sum_{j}\frac{1}{2m}\left[\hat{\boldsymbol{p}}_{j}-{\boldsymbol{A}}(\hat{\boldsymbol{r}}_{j})\right]^{2}+\sum_{j<k}V(\hat{\boldsymbol{r}}_{j}-\hat{\boldsymbol{r}}_{k}),$ (1)
where ${\boldsymbol{A}}(\boldsymbol{r})=m\Omega\hat{z}\times\boldsymbol{r}$ is the vector potential of a uniform magnetic field $\boldsymbol{B}=2m\Omega\hat{z}$ and $V(\boldsymbol{r})$ is the pairwise particle interaction. This Hamiltonian is equivalent to that of a harmonically trapped system in the rotating frame in the rapid rotation limit, i.e., $\hat{H}=\hat{H}^{\prime}-\Omega\hat{L}_{z}$, where $\hat{H}^{\prime}=\sum_{j}(\hat{\boldsymbol{p}}_{j}^{2}/2m+m\Omega^{2}\hat{\boldsymbol{r}}_{j}^{2}/2)+\sum_{j<k}V(\hat{\boldsymbol{r}}_{j}-\hat{\boldsymbol{r}}_{k})$ is the Hamiltonian for the trapped system, $\Omega$ is both the angular velocity of rotation and the trap frequency and $\hat{L}_{z}$ is the total angular momentum. Due to the rotational invariance, an egienstate $|\Psi_{\alpha}\rangle$ of this trapped system $\hat{H}^{\prime}$ is also an egienstate of $\hat{H}$ with energy $E_{\alpha}$. Such a state is naturally localized around the origin (i.e., the center of the trap) as the expectation value of the center of mass coordinate is zero. Let’s consider the dynamical evolution of the system when prepared at $t=0$ in the state
$|\Psi(0)\rangle\equiv e^{i\left({\boldsymbol{p}}_{0}\cdot\hat{\boldsymbol{R}}-\boldsymbol{r}_{0}\cdot\hat{\boldsymbol{P}}\right)}|\Psi_{\alpha}\rangle,$ (2)
where $\hat{\boldsymbol{R}}=\frac{1}{N}\sum_{j}\hat{\boldsymbol{r}}_{j}$ is the center of mass operator and $\hat{\boldsymbol{P}}=\sum_{j}\hat{\boldsymbol{p}}_{j}$ is the total canonical momentum operator. The dynamical state at time $t$ is ($\hbar=1$ throughout the paper)
$|\Psi(t)\rangle=e^{-iE_{\alpha}t}e^{i\left({\boldsymbol{p}}_{0}\cdot\hat{\boldsymbol{R}}(-t)-\boldsymbol{r}_{0}\cdot\hat{\boldsymbol{P}}(-t)\right)}|\Psi_{\alpha}\rangle,$ (3)
where $\hat{\boldsymbol{R}}(t)=e^{i\hat{H}t}\hat{\boldsymbol{R}}e^{-i\hat{H}t}$ and $\hat{\boldsymbol{P}}(t)=e^{i\hat{H}t}\hat{\boldsymbol{P}}e^{-i\hat{H}t}$ are Heisenberg operators. By means of the Heisenberg equation of motion we find [34]
${\boldsymbol{p}}_{0}\cdot\hat{\boldsymbol{R}}(-t)-\boldsymbol{r}_{0}\cdot\hat{\boldsymbol{P}}(-t)={\boldsymbol{p}}_{0}(t)\cdot\hat{\boldsymbol{R}}-\boldsymbol{r}_{0}(t)\cdot\hat{\boldsymbol{P}}$ (4)
such that
$|\Psi(t)\rangle=e^{-iE_{\alpha}t}e^{i\left({\boldsymbol{p}}_{0}(t)\cdot\hat{\boldsymbol{R}}-\boldsymbol{r}_{0}(t)\cdot\hat{\boldsymbol{P}}\right)}|\Psi_{\alpha}\rangle.$ (5)
Here $\boldsymbol{r}_{0}(t)=\boldsymbol{r}_{0}-{\boldsymbol{d}}(0)+{\boldsymbol{d}}(t)$ and ${\boldsymbol{p}}_{0}(t)=M\dot{\boldsymbol{r}}_{0}(t)+M\Omega\hat{z}\times\boldsymbol{r}_{0}(t)$, where ${\boldsymbol{d}}(t)=(d_{x}(t),d_{y}(t))$ is given by
$\begin{pmatrix}d_{x}(t)\\ d_{y}(t)\end{pmatrix}=\begin{pmatrix}\cos\omega_{c}t&\sin\omega_{c}t\\ -\sin\omega_{c}t&\cos\omega_{c}t\end{pmatrix}\begin{pmatrix}d_{x}(0)\\ d_{y}(0)\end{pmatrix},$ (6)
with ${\boldsymbol{d}}(0)=\frac{1}{2}\boldsymbol{r}_{0}+\frac{1}{2M\Omega}\hat{z}\times{\boldsymbol{p}}_{0}$ and $\omega_{c}=2\Omega$.
In general, $\boldsymbol{r}_{0}(t)$ and ${\boldsymbol{p}}_{0}(t)$ are time-dependent functions except when $\boldsymbol{r}_{0}=-\frac{1}{M\Omega}\hat{z}\times{\boldsymbol{p}}_{0}$ for which ${\boldsymbol{d}}(0)=0$. In this case, both $\boldsymbol{r}_{0}(t)=\boldsymbol{r}_{0}$ and ${\boldsymbol{p}}_{0}(t)={\boldsymbol{p}}_{0}$ are time-independent and we obtain the following eigenstates from Eq. (5)
$|\Psi\rangle=e^{-i\left(\hat{\boldsymbol{P}}+M\Omega\hat{z}\times\hat{\boldsymbol{R}}\right)\cdot\boldsymbol{r}_{0}}|\Psi_{\alpha}\rangle.$ (7)
Each of these states shares the same energy $E_{\alpha}$ and is related to the original localized state $|\Psi_{\alpha}\rangle$ by a magnetic translation of the center of mass. More explicitly, the many-body wave function of such a state is
$\Psi(\boldsymbol{r}_{1},\cdots,\boldsymbol{r}_{N})$ $=e^{i\sum_{j}{\boldsymbol{A}}(\boldsymbol{r}_{0})\cdot\boldsymbol{r}_{j}}$
×$\times\Psi_{\alpha}(\boldsymbol{r}_{1}-\boldsymbol{r}_{0},\cdots,\boldsymbol{r}_{N}-\boldsymbol{r}_{0}),$ (8)
where an irrelevant constant phase factor is dropped. When $\boldsymbol{r}_{0}\neq-\frac{1}{M\Omega}\hat{z}\times{\boldsymbol{p}}_{0}$, it is clear that $\boldsymbol{r}_{0}(t)$ and ${\boldsymbol{p}}_{0}(t)$, respectively, represent the trajectory and the canonical momentum of a particle of mass $M=Nm$ performing the cyclotron motion of frequency $\omega_{c}=2\Omega$, whose orbit is centered at $\boldsymbol{r}_{0}$ and has radius given by ${\boldsymbol{d}}(0)$. In this case, the many-body wave function can be written as
$\Psi(\boldsymbol{r}_{1},\cdots,\boldsymbol{r}_{N};t)$ $=e^{-iE_{\alpha}t}e^{i\sum_{j}[m{\dot{\boldsymbol{r}}}_{0}(t)+{\boldsymbol{A}}(\boldsymbol{r}_{0}(t))]\cdot\boldsymbol{r}_{j}}$
×$\times\Psi_{\alpha}(\boldsymbol{r}_{1}-\boldsymbol{r}_{0}(t),\cdots,\boldsymbol{r}_{N}-\boldsymbol{r}_{0}(t)).$ (9)
This means that for ${\boldsymbol{d}}(0)\neq 0$ the probability density of this state $|\Psi(\boldsymbol{r}_{1},\cdots,\boldsymbol{r}_{N};t)|^{2}=|\Psi_{\alpha}(\boldsymbol{r}_{1}-\boldsymbol{r}_{0}(t),\cdots,\boldsymbol{r}_{N}-\boldsymbol{r}_{0}(t))|^{2}$ simply performs cyclotron motion rigidly following the trajectory given by $\boldsymbol{r}_{0}(t)$. Physically our theorem reflects the fact that 2D systems in a magnetic field are similar to those confined in a harmonic potential in that the center of mass degree of freedom is decoupled from the internal ones even in the presence of interaction [35, 36, 37, 38]. We emphasize that the theorem is universal in the sense that it does not depend on particle statistics nor on any specific form of the pairwise interaction. In the following we will make use of this theorem to construct various droplet states for a rapidly rotating BEC.
Single droplet state.—In cold atomic experiments, quasi-2D confinement is ensured by making the axial trap frequency $\omega_{z}$ much larger than the radial frequency $\omega_{\perp}$, such that the effective 2D interaction strength is $g=a_{s}\sqrt{{8\pi\omega_{z}}/{m}}$ where $a_{s}$ is the s-wave scattering length. In the rapid rotation limit $\Omega=\omega_{\perp}$, the quasi-2D Bose condensate in the rotating frame is described by the Gross-Pitaevskii equation [39, 40, 41]
$i\frac{\partial}{\partial t}\psi=\left[-\frac{1}{2m}(\nabla-i{\boldsymbol{A}}(\boldsymbol{r}))^{2}+g|\psi|^{2}\right]\psi,$ (10)
where the condensate wave function $\psi$ is normalized to the number of atoms $N$.
Figure 1: A droplet centered at the origin (left), at $\boldsymbol{r}_{0}=(-7,7)\ell_{B}$ (middle) and performing the cyclotron motion as a whole (right). The arrows denote the current field in the droplets. Here $N=8\times 10^{3}$, $g=1.25\times 10^{-2}\omega_{c}\ell_{B}^{2}$, $\omega_{c}=2\pi\times 177.2$ Hz and $\ell_{B}\approx 1.6\mu$m. The last three are common parameters used in all calculations.
In the first instance, we may look for droplet states from stationary solutions to the above equation. There are two types of stationary solutions that can be easily verified. The first is $\psi(\boldsymbol{r},t)=\psi_{0}(\boldsymbol{r})e^{-i\mu t}$, where $\psi_{0}(\boldsymbol{r})$ is the ground state of the static trapped condensate satisfying
$\left(-\nabla^{2}/2m+m\Omega^{2}r^{2}/2+g|\psi_{0}|^{2}\right)\psi_{0}(r)=\mu\psi_{0}(r),$ (11)
and $\mu$ is the chemical potential. The second is $\psi(\boldsymbol{r},t)=\tilde{\psi}_{0}(x)e^{-im\Omega xy}e^{-i\tilde{\mu}t}$ [42], where $\tilde{\psi}_{0}(x)$ is given by
$\left(-\partial_{x}^{2}/2m+2m\Omega^{2}x^{2}+g|\tilde{\psi}_{0}|^{2}\right)\tilde{\psi}_{0}(x)=\tilde{\mu}\tilde{\psi}_{0}(x).$ (12)
We focus on the former because the second type is known to be dynamically unstable [29]. From Eq. (8) and the relation between the condensate wave function and the many-body wave function, we find that the state
$\psi_{0}(\boldsymbol{r};\boldsymbol{r}_{0})\equiv e^{i{\boldsymbol{A}}(\boldsymbol{r}_{0})\cdot\boldsymbol{r}}\psi_{0}(\boldsymbol{r}-\boldsymbol{r}_{0})$ (13)
must also be a stationary solution to Eq. (10). Aside from these stationary solutions, Eq. (9) suggests that Eq. (10) must accept dynamical states $\psi(\boldsymbol{r},t)=\psi_{0}(\boldsymbol{r};\boldsymbol{r}_{0}(t))e^{-i\mu t}$ as solutions, where
$\psi_{0}(\boldsymbol{r};\boldsymbol{r}_{0}(t))\equiv e^{i[m{\dot{\boldsymbol{r}}}_{0}(t)+{\boldsymbol{A}}(\boldsymbol{r}_{0}(t))]\cdot\boldsymbol{r}}\psi_{0}(\boldsymbol{r}-\boldsymbol{r}_{0}(t))$ (14)
describes a condensate performing the cyclotron motion $\boldsymbol{r}_{0}(t)$ as a whole. Examples for $\psi_{0}(\boldsymbol{r})$ and the related two states $\psi_{0}(\boldsymbol{r};\boldsymbol{r}_{0})$ and $\psi_{0}(\boldsymbol{r};\boldsymbol{r}_{0}(t))$ are shown in Fig. 1. Naturally, the stationary and the moving droplets can be distinguished by the current field ${\boldsymbol{j}}(\boldsymbol{r},t)=m|\psi|^{2}\boldsymbol{v}(\boldsymbol{r},t)$, where
$\boldsymbol{v}(\boldsymbol{r},t)=\mbox{\boldmath$\nabla$}\varphi(\boldsymbol{r},t)-{\boldsymbol{A}}(\boldsymbol{r})$ (15)
is the velocity field. Here $\varphi(\boldsymbol{r},t)$ is the phase of the condensate wave function $\psi(\boldsymbol{r},t)=|\psi(\boldsymbol{r},t)|e^{i\varphi(\boldsymbol{r},t)}$.
Figure 2: (a) Quasi-particle spectrum of a single droplet state. (b) Dynamic evolution of a droplet perturbed from equilibrium. Here $T=2\pi/\Omega$ and $N=800$
Dynamical stability.— For the states in Eq. (13) and Eq. (14) to be genuine droplet states, they must be dynamically stable. Since they are internally the same state as $\psi_{0}(\boldsymbol{r})$, it is sufficient to examine the dynamical stability of the latter. For this purpose, we calculate the quasi-particle spectrum in the rotating frame using the Bogoliubov-de Gennes (BdG) equation, which can be derived by substituting $\psi(\boldsymbol{r},t)=\psi_{0}(\boldsymbol{r})e^{-i\mu t}+\delta\psi(\boldsymbol{r},t)$ in Eq. (10), where
$\delta\psi(\boldsymbol{r},t)=e^{-i\mu t}\sum_{nl}\left[u_{nl}(\boldsymbol{r})e^{-i\epsilon_{nl}t}-v^{*}_{nl}(\boldsymbol{r})e^{i\epsilon_{nl}t}\right].$ (16)
The Bogoliubov amplitudes $u_{nl}(\boldsymbol{r}),v_{nl}(\boldsymbol{r})$ and the excitation energy $\epsilon_{nl}$ are labeled by the principal quantum number $n$ and the angular momentum quantum number $l$. The amplitudes can be further expanded in terms of the Landau level wave functions $\phi_{n^{\prime}l}(\boldsymbol{r})$ as $u_{nl}(\boldsymbol{r})=\sum_{n^{\prime}}U_{nn^{\prime}}\phi_{n^{\prime}l}(\boldsymbol{r})$ and $v_{nl}(\boldsymbol{r})=\sum_{n^{\prime}}V_{nn^{\prime}}\phi_{n^{\prime}l}(\boldsymbol{r})$. In symmetric gauge, the $n^{\prime}$-th Landau level wave function is
$\phi_{n^{\prime}l}(r,\theta)=\mathcal{N}\left(\frac{r}{\sqrt{2}\ell_{B}}\right)^{|l|}L_{n^{\prime}}^{(|l|)}\left(\frac{r^{2}}{2\ell_{B}^{2}}\right)e^{il\theta}e^{-\frac{r^{2}}{4\ell_{B}^{2}}},$ (17)
where $\mathcal{N}=\sqrt{\frac{n^{\prime}!}{2\pi\ell_{B}^{2}(n^{\prime}+|l|)!}}$ is the normalization constant, $\ell_{B}=\sqrt{1/(2m\Omega)}$ is the magnetic length and $L_{n}^{(l)}(x)$ is the generalized Laguerre polynomials $L_{n}^{(l)}(x)=\frac{1}{n!}x^{-l}e^{x}\frac{d^{n}}{dx^{n}}\left(e^{-x}x^{n+l}\right)$. As the angular momentum is a good quantum number, the BdG equation takes the form of
$\sum_{n^{\prime}}\left[\mathcal{L}_{n^{\prime\prime}n^{\prime}}{(l)}U_{nn^{\prime}}-\mathcal{M}_{n^{\prime\prime}n^{\prime}}{(l)}V_{nn^{\prime}}\right]$ $=\epsilon_{nl}U_{nn^{\prime\prime}};$ (18)
$\sum_{n^{\prime}}\left[\mathcal{M}_{n^{\prime\prime}n^{\prime}}{(-l)}U_{nn^{\prime}}-\mathcal{L}_{n^{\prime\prime}n^{\prime}}{(-l)}V_{nn^{\prime}}\right]$ $=\epsilon_{nl}V_{nn^{\prime\prime}},$ (19)
[…]
Crucially, we find that the excitation spectra obtained by solving Eqs. (18)-(19) are all real (see Fig. 2(left)), ensuring that the state $\psi_{0}(\boldsymbol{r})$ is dynamically stable. This is in stark contrast with the excitation spectra of the second type of stationary state $\tilde{\psi}_{0}(x)e^{-im\Omega xy}$ in Eq. (12) which always contains an imaginary branch underlying the exponential instability of the state. The dynamical stability of the droplet is also directly verified by perturbing the condensate from equilibrium [see Fig. 2(right)]. For example, when slightly squeezed along one direction, the droplet performs oscillations that are well accounted for by excitations determined by the BdG equation [34]. Remarkably, even under strong perturbation for which the linear stability analysis is no longer applicable, the droplet is still stable against further dissociating into smaller subsystems or vortex states [34]. Such linear and non-linear stability are essential characteristics of the droplet state.
[…]
$I(t)$ $=({2}/{m\Omega^{2}})(E_{\rm tot}+\Delta E\cos 2\omega t);$ (20)
$\boldsymbol{L}_{\rm p}(t)$ $=\boldsymbol{L}_{\rm c}-I(t),$ (21)
[…]
Figure 3: Stationary droplet arrays. Densities of the droplet arrays at $t=4T$ are shown, which are indistinguishable from those at $t=0$. The parameters here are $N=10^{5}$ (left) and $N=8\times 10^{4}$ (right).
Stationary and rotating droplet arrays.—We now show that the single droplet state in Eq. (13) can be used in constructing stationary droplet arrays of any configuration. Let’s consider a coherent superposition of a number of single droplet states $\psi_{j}$, each containing $N_{j}$ atoms and centered at $\boldsymbol{r}_{j}$, i.e.,
$\psi(\boldsymbol{r},t)=\sum_{j}e^{i{\boldsymbol{A}}(\boldsymbol{r}_{j})\cdot\boldsymbol{r}}\psi_{j}(\boldsymbol{r}-\boldsymbol{r}_{j})e^{-i\mu_{j}t},$ (22)
where $\mu_{j}$ is the chemical potential of the $j$-th droplet. Since each droplet is characterized by a finite radius, we can choose ${\boldsymbol{r}_{j}}$ to be sufficiently apart from each other such that there is no overlap between any two droplets, i.e., $\psi_{i}(\boldsymbol{r}-\boldsymbol{r}_{i})\psi_{j}(\boldsymbol{r}-\boldsymbol{r}_{j})=0$ for any $i\neq j$. In this case, it can be easily verified $\psi(\boldsymbol{r},t)$ in Eq. (22) is a solution to Eq. (10). Since there is no overlap between different droplets, the total density of the condensate is stationary, i.e.,
$\rho(\boldsymbol{r})=\sum_{j}|\psi_{j}(\boldsymbol{r}-\boldsymbol{r}_{j})|^{2}.$ (23)
We emphasize that such a droplet array can be of any configuration as long as different droplets have no overlaps. As two simple examples, we show in Fig. 3 a 2D droplet lattice and a circular droplet array constructed using single droplet states containing the same number of atoms. By time evolving these states we have verified that the densities of the arrays so constructed are indeed stationary. Similar to our previous discussions of the moving single droplet we can also construct droplet arrays that performs rigid-body rotations at the cyclotron frequency $\omega_{c}$ by taking
$\psi=\sum_{j}e^{i[m\dot{\boldsymbol{r}}_{j}(t)+{\boldsymbol{A}}(\boldsymbol{r}_{j}(t))]\cdot\boldsymbol{r}}\psi_{j}(\boldsymbol{r}-\boldsymbol{r}_{j}(t))e^{-i\mu_{j}t},$ (24)
where $\boldsymbol{r}_{j}(t)$ are cyclotron trajectories all centered at the origin.
Formation of dynamic droplet arrays.—
Figure 4: Formation of dynamic droplet arrays. A uniform condensate (upper left) with phase designed according to Eq. (25) evolves into a droplet lattice (lower left). Here $N=2.4\times 10^{5}$ and the dimension of the initial condensate is $24\sqrt{2}\ell_{B}\times 80\sqrt{2}\ell_{B}$. A ring shaped uniform condensate (upper right) with phase designed according to Eq. (26) evolves into a circular droplet array (lower right). Here $N=8\times 10^{4}$ and the outer radius and the width of initial condensate are $14\sqrt{2}\ell_{B}$ and $4\sqrt{2}\ell_{B}$ respectively.
After establishing various droplet array solutions to the time-dependent GP equation of the rapidly rotating BEC, a natural question concerns the stability of these states. This is a considerably more complex problem than the stability of a single droplet because under external perturbations the interaction between adjacent droplets will come into play. Rather than calculating the excitation spectrum, here we demonstrate that droplet arrays can emerge spontaneously out of initial states that are significantly deviated from the stationary droplet array states. The droplet arrays so obtained are necessarily dynamic and can be viewed as the excited states of the stationary droplet array; their formation is alternative evidence that droplet arrays are indeed stable structures of the rapidly rotating BEC.
Figure 5: The SO(2,1) dynamics of one of the droplets in the circular droplet array (the one encircled by dashed lines in Fig. 4). The left figure displays the temporal evolution of $I(t)$ and $\boldsymbol{L}_{\rm p}(t)$, both calculated relative to center of the droplet. The right shows the snapshots of the droplet at various times.
Figure 6: A uniform strip of condensate with phase designed according to Eq. (27) evolves into a droplet array rotating at the cyclotron frequency $\omega_{c}=2\Omega$. Here the dimension of the initial condensate is $4\sqrt{2}\ell_{B}\times 80\sqrt{2}\ell_{B}$ and $N=8\times 10^{4}$.
In fact, it has been shown both theoretically and experimentally [29] that a 1D droplet array can be formed by preparing the system in the stationary state $\psi(\boldsymbol{r})=\tilde{\psi}_{0}(x)e^{-im\Omega xy}$ where $\tilde{\psi}_{0}(x)$ is determined by Eq. (12). Interestingly we find that the phase $\varphi=-m\Omega xy$ of the initial wave function, which gives rise to a sheared velocity profile, always results in the formation of the 1D droplet array regardless of the initial density profile. Motivated by this insight, we find that it is possible to engineer the phase distribution of the initial wave function so as to achieve sheared velocity profiles conducive to the formation of 2D droplet arrays. For instance, a 2D droplet array with the lattice configuration can be obtained by the following initial phase
$\varphi(x,y)=-m\Omega(x-2jl_{D})y$ (25)
for $(j-1)l_{D}/2\leq x\leq(j+1)l_{D}/2$, where $l_{D}$ is a parameter that determines the size of the eventual droplets and $j=0,\pm 1,\pm 2,\cdots$. For the circular droplet array, the suitable initial phase (in polar coordinates) is
$\varphi(r,\theta)=m\Omega r_{0}^{2}\theta,$ (26)
for $0<\theta\leq 2\pi$ and $r_{\rm inner}\leq r\leq r_{\rm outer}$ , where $r_{0}=(r_{\rm inner}+r_{\rm outer})/2$. The corresponding droplet arrays resulted from these two phase distributions are shown in Fig. 4. As mentioned earlier, once the droplet arrays are formed they are still in a dynamic state, as each of its droplets is in a constant, combined motion of stretching and rotation [34]. To make contact with previously discussed single droplet dynamics, we calculate the moment of inertia and the physical angular momentum for a droplet in the circular array, which shows perfect agreements with Eqs. (20)-(21) (see Fig. 5). As a final example of the phase engineering approach to the formation of droplet arrays, we find that a 1D droplet array rotating at the cyclotron frequency $\omega_{c}$ can be achieved by the initial phase
$\varphi(x,y)=m\Omega xy,$ (27)
which generates a velocity field ${\bf v}=(2\Omega y,0)$ as shown in Fig. 6. The rotation of the array can be understood from Eq. (24) if we neglect the internal dynamics of individual droplets.
Conclusions.—In summary, we have systematically studied the formation and dynamics of single droplet and droplet arrays in rapidly rotating BECs. Our theorem of localized many-body states of 2D interacting systems in a magnetic field provides a fundamental understanding for why droplet states exist in these rotating BECs. Our extensive numerical simulations provide further insight into the dynamical properties of the single droplet and droplet arrays. In particular, we have demonstrated the spontaneous formation of various droplet arrays out of phase-engineered initial states. Such an approach may be adopted in experiments and pave the way for the future study of novel dynamics in rotating BEC.
Acknowledgements. We thank Yueheng Lan and Ping Fang for helpful discussions. This work is supported by NSFC under Grant No. 12374243 (MYS) and Xiaomi Young Talents Program. Z.W. is supported by NSFC (Grant No. 12474264), Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2404007) and National Key R&D Program of China (Grant No. 2022YFA1404103).